A wedge shaped film is obtained by placing a thin wire between two plane glass  plates at one end keeping them in contact with each other at the other end. When  the film is illuminated by light of wavelength 6000 Å, 40 fringes are observed.  Calculate the radius of the wire.


1
Expert's answer
2021-06-07T09:29:25-0400

"r=\frac{\lambda m}{2}=\frac{6\cdot 10^{-7}\cdot 40}{2}=12~\mu m."